# Program Synthesis Dialog Agents for Interactive Decision-Making

### v2

@@ Abstract
Many real-world eligibility problems, ranging from medical diagnosis to tax planning, can be mapped to decision problems expressed in natural language, wherein a model must make a binary choice based on the features of the user. Large-scale domains such as legal codes or frequently updated funding opportunities render human annotation (e.g., web forms or decision trees) impractical, suggesting a need for agents that can automatically assist in decision-making. Since relevant information is often only known to the user, it is important that these agents can ask the right questions. Since agents determine when to terminate a conversation, they face a trade-off between accuracy and total questions asked, a useful metric for both user experience and cost. To evaluate this task, we propose BeNYfits, a new benchmark for determining user eligibility for multiple overlapping social benefits opportunities through interactive decision-making. Our experiments show that current language models struggle with frequent hallucinations, with GPT-4o scoring only 35.7 F1 using a ReAct-style chain-of-thought. We therefore introduce ProADA, a novel approach that uses program synthesis to assist in decision-making by mapping dialog planning to a code generation problem and using gaps in structured data to determine the best next action. Our agent, ProADA, improves the F1 score to 55.6 while using nearly the same number of dialog turns.
@@
Matthew Toles1, Nikhil Balwani1,2*, Rattandeep Singh1, Valentina Giulia Sartori Rodriguez1,3, Zhou Yu1 1Columbia University, 2Amazon, 3Sciences Po Paris Correspondence: m.toles@columbia.edu This work was finished before author joined Amazon
@@ 1 Introduction
The improved capabilities of large language models have refocused attention away from traditional benchmarks and towards real-world tasks where automated systems could broadly benefit the public, such as improving access to public services. Many such opportunities require determining the user’s eligibility based on the user’s features and the task at hand, formally referred to as decision problems. In adaptive decision scenarios, where information is revealed iteratively (e.g., medical diagnosis), one also wishes to minimize the number of queries. For many important decision problems, the requirements are written in natural language, which must be translated into formal language or interpreted logically. Also, critical information may be user-specific and known only to certain people, meaning that it often must be requested through dialog. Finally, the diversity of problems in the real world places a premium on whether agents can generalize information gathering and logical reasoning to new domains.
@@
User-facing decision problems have traditionally been solved using hard-coded forms (as in the American tax filing software TurboTax [^1] https://turbotax.intuit.com/) or dialog trees (as in video games). However, hard-coded solutions struggle to generalize or extend to web-scale decision problems; opportunities found by web-scraping, crowd-sourcing, or from extremely large corpora such as national tax codes may be challenging to formalize, let alone decide on in real-time. Methods to approach adaptive decision problems include multi-armed bandits, reinforcement learning, dynamic programming, and decision trees, but adapting these strategies to online natural language tasks is not trivial. Recently, large language models have improved on a wide range of related tasks. However, language models are known to struggle with reasoning over long contexts and hallucinating unstated information, both of which, as we will see, are common in adaptive decision problems.
@@
To evaluate interactive decision-making, in Section 2 we introduce BeNYfits, a language model agent task for determining user eligibility for real-world public benefits opportunities with overlapping eligibility requirements. 1. In the single-opportunity scenario, the assistant agent could simply repeat the requirements and ask the user if they qualify. However, BeNYfits’ overlapping requirements present an interesting optimization challenge for dialog planning: how should models “merge" eligibility requirements to avoid duplicating questions and maximize information gain? We find that current large language models, including GPT-4o, struggle to perform significantly better than chance at determining user eligibility, suffering from hallucination, poor reasoning under uncertainty, overconfidence, and lost-in-the-middle problems observed in prior work Huang et al. (2024).
@@
Given these weaknesses, we introduce a method for an agent that, given a natural language description of the user-facing decision problem, generates a program to request user input conversationally to solve the problem. Specifically, we construct an agent consisting of a code module, which conducts dialog planning in the form of a Python program, and a dialog module, which asks questions based on the program state. The agent then uses the dialog module to parse the user’s response into structured data. Our approach exploits a distinction that we observe between how dialog and code generation models handle long-range planning and uncertainty. Effectively, the formality and long-range conditional dependencies present in program synthesis training make them strong candidates for dialog planning. As a key contribution, in Section 3 we present Program Synthesis Adaptive Decision Agent, or ProADA, an agent that, given a natural language policy of a decision problem, generates Python code to structure the decision-making process and request minimal user input to make the correct decision.
@@
2. A general and effective agent for adaptive decision-making in dialog that exploits program synthesis and tool use to plan dialog and adaptively request user information, improving both F1 score and dialog completion speed.  [^2] https://github.com/mtoles/BeNYfits-ProADA/
@@ 2 BeNYfits: An Agent Benchmark for Public Benefit Eligibility Decisions
The determination of eligibility for many public opportunities, such as tax credits, scholarships, research funding opportunities, business incentives, charities, job listings, and social services, can be reduced to a binary decision problem. Since many requirements, such as age and income, overlap between programs, this creates an opportunity for agent assistants to make more efficient and adaptive decisions as compared to traditional methods like static web forms. At the same time, determinations often require domain-specific knowledge to make accurate determinations, presenting a challenge in natural language understanding. We present BeNYfits, a benchmark for decision-making on public benefits eligibility. In BeNYfits, the agent’s goal is to help users navigate complex decision-making processes, focusing on objective, documented opportunities requiring user input, then make a final determination on the user’s eligibility based on the dialog, in the minimum number of dialog turns.
@@ 2.1 Efficiency, Generalization, and User Experience
Traditional methods for determining eligibility present several opportunities for improvement by intelligent agent assistants. For small numbers of opportunities, we might convert natural language requirements into a web form or static chatbot dialog tree serving as an interface for hard-coded checking logic, similar to TurboTax. However, this approach has several drawbacks. First, many eligibility requirements are updated without notice, often annually, meaning eligibility performance will degrade over time without ongoing maintenance. Second, opportunities may be crowdsourced or scraped from the Web dynamically, rendering manual coding impractical in favor of more generalizable, lower-latency solutions, such as language agents. Moreover, due to the specificity of eligibility requirements, a potential applicant may need to find and examine many opportunities before discovering one for which they qualify. Furthermore, requirements between similar opportunities frequently overlap, forcing users to answer the same questions repeatedly. On the other hand, users may waste a lot of time if they discover that they are ineligible very late in the examination process, presenting an opportunity for adaptive decision-making algorithms. An intelligent agent, however, should adaptively query the user for only the minimum necessary information, saving time, and improving user experience.
@@ 2.2 Opportunity Requirements
We source the plain English eligibility requirements for 82 benefits opportunities from NYC Open Data [^3] https://data.cityofnewyork.us/. We minimally edit requirements manually to remove ambiguity ("may be eligible"), future expectations ("can commit to"), and dates ("since 2023"). Opportunities include tax credits, youth programming, housing, nutrition assistance, healthcare, parental services, and career advancement, among other categories. Eligibility requirements range from broad (State ID card: all residents age 10+) to extremely specific (Air conditioner subsidy: ten independent requirements). Opportunities may apply to either the individual or the household as a whole, offering additional logical complexity. Opportunities depend on 1-18 unique user features each (mean: 4.66 standard deviation: 3.56). Each user feature appears in 1-52 opportunities (mean: 3.25, standard deviation: 6.66), falling on a long-tailed distribution (Figure 2). We define a household as eligible for an opportunity if any of its members are eligible.
@@ 2.5 Diverse Dataset
Due to the time and cost of benchmarking new models, we attempt to construct the smallest possible dataset with the most "coverage" of unique traces to qualification or disqualification using input fuzzing. Each example consists of a simulated user, a subset of opportunities in which they are interested, and the ground truth eligibility for those opportunities. We first randomly generate 10,000 simulated users, sampling each variable from a distribution chosen to produce a balanced distribution across traces, typically uniformly, except when eligibility is based on thresholds of numeric features. We then greedily add households to the dataset whose trace through all eligibility checkers contributes the most lines not yet present in the dataset. Finally, we greedily remove opportunities from simulated users if the user-opportunity trace contributes no unique lines. We refer to this dataset as the Diverse Dataset, which contains only 56 of the original 10,000 households (305 of 82,000 user-opportunity pairs) but covers the same traces as the full set. Each household is interested in between 1-10 opportunities, with a mean of 5.4.
@@ 2.7 Dialog Loop
Agents are provided eligibility requirements and then must determine simulated user eligibility by asking questions, one at a time. After each response, the agent is prompted with Ready, where it is asked if it has enough information to determine the eligibility of the user with certainty (Figure 1). If the agent responds with True, it is prompted to Predict the user’s eligibility for each opportunity. Otherwise, it asks another question. We limit conversations to 20 questions per opportunity, to a maximum of 100 questions.
@@ 3 ProADA
To best explain ProADA, we instantiate it in the context of our proposed BeNYfits benchmark. Before starting a dialog, ProADA uses the code generation model to convert the eligibility requirements in natural language into a Python Decide Checker tool used by the agent 3. Decide is a Python function that takes a UserFeatures dictionary containing known user properties (e.g., ‘‘homeless_or_runaway") as input and outputs the household eligibility. For each key used to access UserFeatures, the code generation model defines a type (int, float, etc.) constraint, or a list of string choices that the feature can take. At the start of the dialog, the agent runs the Decide tool, passing in an empty UserFeatures dictionary, since it knows nothing about the user yet. An empty dictionary would normally cause a key error, which we exploit by wrapping Decide in a try/catch block. In an exception, the agent passes the offending key, relevant code, eligibility requirements, and dialog history to the dialog module. The dialog module constructs a question seeking the necessary information (“Are you a homeless or a runaway youth?") and presents it to the user. The dialog module then converts the user’s response (“I am a homeless youth") into a valid value according to the predetermined constraint using constrained generation, storing the key-value pair (‘‘homeless_or_runaway": ‘‘yes") in UserFeatures. The agent repeats running Decide with the updated UserFeatures dictionary until it returns a value.
@@ 4 Experimental Setup
As baselines, we choose Llama 3.1 Instruct 8B and 70B, as well as GPT-4o. In direct prompting, we instruct these models to assess readiness and generate questions at each step in the dialog loop, and finally to predict eligibility. We also assess prompting models to conduct ReAct-style chain-of-thought before each step. During Readly and Decide, we use constrained decoding to ensure ProADA and baseline models generate a valid output. For ProADA, we choose the same models for the dialog module and always use GPT-4o for the code generation module. We choose Llama 3.1 Instruct 70B to implement our simulated user for all experiments, in order to reduce the hallucinations that we observed in smaller 8B parameter models. To reduce memory usage, we use 4-bit quantization on all 70B-parameter models. We report three trials for all tests, except for the following, which we run once, due to resource constraints: GPT-4o + direct prompting, ReAct, and ProADA; and Llama 3.1 70B + ReAct.
To measure human performance on this task, one expert author performed the role of the agent on 39 user-opportunity pairs from the Diverse Dataset, achieving 89.7% accuracy. Upon review, we find all inaccuracies are due to human error rather than unfaithful simulated user responses, suggesting a performance ceiling of nearly 100% on this benchmark.
@@ 5 Experimental Results
Figure 4: Average of Representative and Diverse dataset F1 vs. dialog turns to completion for ProADA and baseline models.
@@
Since our datasets are unbalanced (Diverse: 47.9% positive, Representative: 15.5%), we choose micro F1 as our primary accuracy metric. Let F1 and T be the average F1 score and the number of turns across both datasets, respectively. To reward efficient questioning, we define a turn-weighted F1 score as:
[…]
since dialogs can have at most 100 turns.
@@
We find that our method, ProADA, outperforms all others by a significant margin with a turn-weighted F1 score of 47.7. The turn-weighted F1 score drops to 38.5 when using Llama 3.1 8B versus GPT-4o as the dialog model. However, Llama 3.1 8B + ProADA still surpasses the next best strategy, Llama 3.1 70B + ReAct at 34.1 and GPT-4o + direct prompting, despite using a dialog model with many times fewer parameters. ProADA F1 scores exceed those of equivalent models using direct or ReAct prompting in all but two cases in both datasets (Table 1). Among directly prompted models, GPT-4o and Llama 3.1 8B score highest (29.9 and 31.7, respectively). GPT-4o achieves a relatively high average F1 score (40.8), but uses 27.2 turns per dialog, while Llama 3.1 8B appears to terminate prematurely after only 6.0 turns, achieving only 33.6 F1. Dialog completion speed varies widely across models and strategies, with Llama 3.1 70B + direct prompting frequently hitting the turn limit without terminating, resulting in an average turn count of 81.7. Program synthesis guidance, and to a lesser degree ReAct prompting, appear to moderate the number of turns needed without negatively impacting F1 score.
Dialog Model Diverse Set Representative Set Average Turn-Weighted
Strategy Llama F1 ↑ Turns ↓ F1 ↑ Turns ↓ F1 ↑ Turns ↓ F1 ↑
Random P(True)=0.5 23.6 0.0 48.9 0.0 36.3 0.0 36.3
Direct Prompting Llama 3.1 8B 44.6 5.5 22.5 6.4 33.6 6.0 31.7
Llama 3.1 70B 54.3 41.6 27.2 81.7 40.8 61.7 25.2
GPT-4o 31.2 16.8 45.0 37.6 38.1 27.2 29.9
ReAct Agent Llama 3.1 8B 40.4 9.5 26.1 10.9 33.2 10.2 30.2
Llama 3.1 70B 58.3 9.5 26.1 37.9 42.2 23.7 34.1
GPT-4o 50.4 18.4 20.9 13.3 35.7 15.8 30.8
ProADA (Ours) Llama 3.1 8B 50.6 16.8 41.9 23.4 46.2 20.1 38.5
Llama 3.1 70B 54.0 18.2 49.7 19.8 51.8 19.0 43.5
GPT-4o 50.4 15.7 60.8 17.2 55.6 16.5 47.7

### v3

@@ Abstract
Many real-world eligibility problems, ranging from medical diagnosis to tax planning, can be mapped to decision problems expressed in natural language, wherein a model must make a binary choice based on the features of the user. Large-scale domains such as legal codes or frequently updated funding opportunities render human annotation (e.g., web forms or decision trees) impractical, suggesting a need for agents that can automatically assist in decision-making. Since relevant information is often only known to the user, it is important that these agents can ask the right questions. To evaluate this task, we propose BeNYfits, a new benchmark for determining user eligibility for multiple overlapping social benefits opportunities through interactive decision-making. Our experiments show that current language models struggle with frequent hallucinations, with GPT-4o scoring only 35.7 F1 using a ReAct-style chain-of-thought. We therefore introduce ProADA, a novel approach that uses program synthesis to assist in decision-making by mapping dialog planning to a code generation problem and using gaps in structured data to determine the best next action. Our agent, ProADA, improves the F1 score to 56.2 while using nearly the same number of dialog turns. * This work completed before joining Amazon
@@
Matthew Toles1, Nikhil Balwani1,2*, Rattandeep Singh1, Valentina Giulia Sartori Rodriguez1,3, Zhou Yu1 1Columbia University, 2Amazon, 3Sciences Po Paris Correspondence: m.toles@columbia.edu
@@ 1 Introduction
The improved capabilities of large language models have refocused attention away from traditional benchmarks and towards real-world tasks where automated systems could broadly benefit the public, such as improving access to public services. Many such opportunities require determining the user’s eligibility based on the user’s features and the task at hand, formally referred to as decision problems. In adaptive decision scenarios, where information is revealed iteratively (e.g., medical diagnosis), one also wishes to minimize the number of queries. Also, critical information may be user-specific and known only to certain people, meaning that it often must be requested through dialog. Finally, the diversity of problems in the real world places a premium on whether agents can generalize information gathering and logical reasoning to new domains.
@@
User-facing decision problems have traditionally been solved using hard-coded forms (as in the American tax filing software TurboTax [^2] https://turbotax.intuit.com/) or dialog trees (as in video games). However, hard-coded solutions struggle to generalize or extend to web-scale decision problems; opportunities found by web-scraping, crowd-sourcing, or from extremely large corpora such as national tax codes may be challenging to formalize, let alone decide on in real-time. Methods to approach adaptive decision problems include multi-armed bandits, reinforcement learning, dynamic programming, and decision trees, but adapting these strategies to online natural language tasks is not trivial. Recently, large language models have improved on a wide range of related tasks. However, they are known to struggle with reasoning over long contexts and hallucinate unstated information.
@@
To evaluate interactive decision-making, in Section 2 we introduce BeNYfits, a language model agent task for determining user eligibility for real-world public benefits opportunities with overlapping eligibility requirements (Figure 1). In the single-opportunity scenario, the assistant agent could simply repeat the requirements and ask the user if they qualify. However, BeNYfits’ overlapping requirements present an interesting optimization challenge for dialog planning: how should models “merge" eligibility requirements to avoid duplicating questions and maximize information gain? We find that current large language models, including GPT-4o, struggle to perform significantly better than chance at determining user eligibility, suffering from hallucination, poor reasoning under uncertainty, overconfidence, and lost-in-the-middle problems observed in prior work Huang et al. (2024).
@@
Given these weaknesses, we introduce a method for an agent that, given a natural language description of the user-facing decision problem, generates a program to request user input conversationally to solve the problem. Specifically, we construct an agent consisting of a code module, which conducts dialog planning in the form of a Python program, and a dialog module, which asks questions based on the program state. The agent then uses the dialog module to parse the user’s response into structured data. This approach exploits code generation models’ long-range planning and and uncertainty handling to improve task-oriented dialog, as compared to conventional dialog models. As a key contribution, in Section 3 we present Program Synthesis Adaptive Decision Agent, or ProADA, an agent that, given a natural language policy of a decision problem, generates Python code to structure the decision-making process and request minimal user input to make the correct decision.
@@
2. A general and effective agent for adaptive decision-making in dialog that exploits program synthesis and tool use to plan dialog and adaptively request user information, improving both F1 score and dialog completion speed. We will release the model code and maintain a benchmark after publication.
@@ 2 BeNYfits: An Agent Benchmark for Public Benefit Eligibility Decisions
The determination of eligibility for many public opportunities, such as tax credits, scholarships, research funding opportunities, business incentives, charities, job listings, and social services, can be reduced to a binary decision problem. Since many requirements, such as age and income, overlap between programs, this creates an opportunity for agent assistants to make more efficient and adaptive decisions as compared to traditional methods like static web forms. At the same time, determinations often require domain-specific knowledge to make accurate determinations, presenting a challenge in natural language understanding. We present BeNYfits, a benchmark for decision-making on public benefits eligibility. In BeNYfits, the agent’s goal is to help users navigate complex decision-making processes and make a final determination on the user’s eligibility based on the dialog in the minimum number of dialog turns.
@@ 2.1 Efficiency, Generalization, and User Experience
Traditional methods for determining eligibility present several opportunities for improvement by intelligent agent assistants. For small numbers of opportunities, we might convert natural language requirements into a web form or static chatbot dialog tree serving as an interface for hard-coded checking logic, similar to TurboTax. However, this approach has several drawbacks. First, many eligibility requirements are updated without notice, often annually, meaning eligibility performance will degrade over time without ongoing maintenance. Second, opportunities may be crowdsourced or scraped from the Web dynamically, rendering manual coding impractical in favor of more generalizable, lower-latency solutions, such as language agents. Furthermore, requirements between similar opportunities frequently overlap, forcing users to answer the same questions repeatedly. On the other hand, users may waste a lot of time if they discover that they are ineligible very late in the examination process, presenting an opportunity for adaptive decision-making algorithms. An intelligent agent, however, should adaptively query the user for only the minimum necessary information, saving time, and improving user experience.
@@ 2.2 Opportunity Requirements
We source the plain English eligibility requirements for 82 benefits opportunities from NYC Open Data [^3] https://data.cityofnewyork.us/. We minimally edit requirements to remove ambiguity ("may be eligible"), future expectations ("can commit to"), and dates ("since 2023"). Opportunities include tax credits, youth programming, housing, nutrition assistance, healthcare, parental services, and career advancement, among other categories. Eligibility requirements range from broad (State ID card: all residents age 10+) to extremely specific (Air conditioner subsidy: ten independent requirements). Opportunities may apply to either the individual or the household as a whole, offering additional logical complexity. Opportunities depend on 1-18 unique user features each (mean: 4.66 standard deviation: 3.56). Each user feature appears in 1-52 opportunities (mean: 3.25, standard deviation: 6.66), falling on a long-tailed distribution (Figure 2). We define a household as eligible for an opportunity if any of its members are eligible.
@@ 2.5 Diverse Dataset
Due to the time and cost of benchmarking new models, we attempt to construct the smallest possible dataset with the most "coverage" of unique traces to qualification or disqualification using input fuzzing. Each example consists of a simulated user, a subset of opportunities in which they are interested, and the ground truth eligibility for those opportunities. We first randomly generate 10,000 simulated users, sampling each variable from a distribution chosen to produce a balanced sampling across traces, typically uniformly, except when eligibility is based on thresholds of numeric features. We then greedily add households to the dataset whose trace through all eligibility checkers execute the most unique lines of code not yet present in the dataset. Finally, we greedily remove opportunities from simulated users if the user-opportunity trace contributes no unique lines. We refer to this dataset as the Diverse Dataset, which contains only 56 of the original 10,000 households (305 of 82,000 user-opportunity pairs) but covers the same traces as the full set. Each household is interested in between 1-10 opportunities, with a mean of 5.4.
@@ 2.7 Dialog Loop
Agents are provided eligibility requirements and then must determine simulated user eligibility by asking questions, one at a time. After each response, the agent is prompted with Ready, where it is asked if it has enough information to determine the eligibility of the user with certainty (Figure 1). If the agent responds with True, it is prompted to Predict the user’s eligibility for each opportunity. Otherwise, it asks another question. We limit conversations to average 20 questions per opportunity, to a maximum of 100 questions.
@@ 3 ProADA
To best explain ProADA, we instantiate it in the context of our proposed BeNYfits benchmark. Before starting a dialog, ProADA uses the code generation model to convert the eligibility requirements in natural language into a Python Decide Checker tool used by the agent (Figure 3). Decide is a Python function that takes a UserFeatures dictionary containing known user properties (e.g., ‘‘homeless_or_runaway") as input and outputs the household eligibility. For each key used to access UserFeatures, the code generation model defines a type (int, float, etc.) constraint, or a list of string choices that the feature can take. At the start of the dialog, the agent runs the Decide tool, passing in an empty UserFeatures dictionary, since it knows nothing about the user yet. An empty dictionary would normally cause a key error, which we exploit by wrapping Decide in a try/except block. In an exception, the agent passes the offending key, relevant code, eligibility requirements, and dialog history to the dialog module. The dialog module constructs a question seeking the necessary information (“Are you a homeless or a runaway youth?") and presents it to the user. The dialog module then converts the user’s response (“I am a homeless youth") into a valid value according to the predetermined constraint using constrained generation, storing the key-value pair (‘‘homeless_or_runaway": ‘‘yes") in UserFeatures. In the case that the user’s response cannot be mapped to a valid value, ProADA will invoke the Clarity module, which will spend up to three turns attempting to clarify the user’s response (additional details in Appendix C.12). This module allows ProADA to recover from unhappy dialog paths and incoherent user input. Finally, the agent repeats running Decide with the updated UserFeatures dictionary until it returns a value.
@@ 4 Experimental Setup
As baselines, we choose Llama 3.1 Instruct 8B and 70B, as well as GPT-4o. In direct prompting, we instruct these models to assess readiness and generate questions at each step in the dialog loop, and finally to predict eligibility. We also assess prompting models to conduct ReAct-style chain-of-thought before each step Yao et al. (2023). During Readly and Decide, we use constrained decoding to ensure ProADA and baseline models generate a valid output. For ProADA, we choose the same models for the dialog module and always use GPT-4o for the code generation module. We choose Llama 3.1 Instruct 70B to implement our simulated user for all experiments, in order to reduce the hallucinations that we observed in smaller 8B parameter models. To reduce memory usage, we use 4-bit quantization on all 70B-parameter models. We report three trials for Llama 8B, Llama 70B ReAct, and ProADA models, and one trial for others due to resource constraints. To measure human performance on this task, two expert authors performed the role of the agent on 83 user-opportunity pairs from the Diverse Dataset, achieving 85.6% accuracy. Upon review, we find all inaccuracies are due to human error rather than unfaithful simulated user responses, suggesting a performance ceiling near 100% on this benchmark.
@@ 5 Experimental Results
Figure 4: Average of Representative and Diverse dataset F1 vs. dialog turns to completion for ProADA and baseline models. Legend follows the format Strategy (Code Model) - Dialog Model.
@@
Because our datasets are unbalanced (Diverse: 47.9% positive, Representative: 15.5%), we choose micro F1 as our primary accuracy metric. Let F1 and $T$ be the average F1 score and the number of turns across both datasets, respectively. To reward efficient questioning, we define a turn-weighted F1 score as:
$\text{Turn-Weighted F1}=\frac{100\cdot\text{F1}}{T/100+1}$ (1)
We choose this metric because it increases monotonically from 0 to 1 with respect to F1 and decreases monotonically with respect to turns. The inclusion of the +1+1 term prevents small changes in $T$ from drastically impacting turn-weighted F2 when $T$ is small. Holding F1 constant, turn-weighted F1 will range from F1 to $0.5\cdot\text{F1}$ as $T$ increases. Because $T\leq 100$, we normalize it by 100.
@@
We find that our method, ProADA, outperforms all others by a significant margin with a turn-weighted F1 score of 46.7 (56.2 F1, 20.4 turns, Figure 4). The turn-weighted F1 score drops to 43.5 without the Clarity module. However, ProADA still surpasses the next best strategy, Llama 3.1 70B + ReAct (34.1) and GPT-4o + direct prompting (29.9), despite using a dialog model with many times fewer parameters. GPT-4o achieves a relatively high average F1 score (40.8), but uses 27.2 turns per dialog, while Llama 3.1 8B appears to terminate prematurely after only 6.0 turns, achieving only 33.6 F1. Claude 3.7 slightly outperforms GPT-4o when serving as the ProADA code module (47.1 vs 46.7 Turn-Weighted F1), likely due to its internal inference-time scaling. However, GPT-4o outperforms in the direct prompting and ReAct conditions, leading us to choose it as our primary baseline. Dialog completion speed varies widely across models and strategies, with Llama 3.1 70B + direct prompting frequently hitting the turn limit without terminating, resulting in an average turn count of 81.7. Program synthesis guidance, and to a lesser degree ReAct prompting, appear to moderate the number of turns needed without negatively impacting F1 score. We see significantly higher accuracy on the Diverse Dataset compared to the Representative Dataset, possibly because its examples contain roughly half as many opportunities per user.
5.1 User Study
We conducted a user survey n=58 interactions. Each user selected 5 opportunities in which they are personally interested. Half of participants were assigned each to use either ProADA or GPT-4o + ReAct. Users then rated the accuracy and subjective usability on a 5-point likert scale, in accordance with the NASA Task Load Index Hart (2006). We find that ProADA achieves better usability scores in metrics as well as significantly higher accuracy and program discovery with real users (Table 6.2).
Model Diverse Representative Average Weighted
Strategy Dialog Code F1 ↑ Turns ↓ F1 ↑ Turns ↓ F1 ↑ Turns ↓ F1 ↑
Direct Prompting Llama 3.1 8B GPT-4o 44.6 5.5 22.5 6.4 33.6 6.0 31.7
Llama 3.1 70B GPT-4o 54.3 41.6 27.2 81.7 40.8 61.7 25.2
GPT-4o GPT-4o 31.2 16.8 45.0 37.6 38.1 27.2 29.9
GPT-o3 mini GPT-4o 52.9 20.3 15.6 31.2 34.2 25.8 27.2
Claude 3.7 GPT-4o 43.7 11.9 17.9 8.9 30.8 10.4 27.9
ReAct Agent Llama 3.1 8B GPT-4o 40.4 9.5 26.1 10.9 33.2 10.2 30.2
Llama 3.1 70B GPT-4o 58.3 9.5 26.1 37.9 42.2 23.7 34.1
GPT-4o GPT-4o 50.4 18.4 20.9 13.3 35.7 15.8 30.8
Claude 3.7 GPT-4o 41.3 38.7 16.4 28.4 28.8 33.6 21.6
Random P(True=0.5) - 23.6 0 48.9 0 36.3 0 36.3
ProADA (ours) Llama 3.1 70B GPT-4o 58.7 18.8 53.7 22.1 56.2 20.4 46.7
Llama 3.1 70B Claude 3.7 57.0 21.5 58.0 22.4 57.5 22.0 47.1
Llama 3.1 70B Qwen 32B 21.9 10.1 17.9 11.5 19.9 10.8 17.9
ProADA w/o Clarity Llama 3.1 70B GPT-4o 54.0 18.2 49.7 19.8 51.8 19.0 43.5